If you draw a card with a value of five or less from a standard deck of cards, I will pay you $7. If not, you pay me $7. (Aces are considered the highest card in the deck.) Step 2 of 2 : If you played this game 543 times how much would you expect to win or lose? Round your answer to two decimal places. Losses must be entered as negative.
Answer #1

In a standard deck of 52 cards, 16 cards are five or less.

P(five or less) = 16 / 52

E(X) = 7 * 16/52 - 7 * ( 1 - 16/52)

= -2.69

If game played for 543 times,

Expected value = 543 * ( -2.69)

= -1460.67